OpenStudy (xoxolove456):

Please help me with this word problem. Alex can paint a room alone in 8 hours. Working together Alex and Lisa can paint the room in 3 hours. How pin would it take lisa to paint the room alone??????

OpenStudy (xoxolove456):

Long not pin

OpenStudy (luffingsails):

\[1/8 + 1/x = 1/3\] \[1/x = 1/3-1/8\] \[1/x=8/24-3/24\] \[1/x = 5/24\] \[x = 24/5\] So, it takes Lisa 4 and 4/5 hours to paint the room alone.

OpenStudy (luffingsails):

or 4.8 hours if you wish.

OpenStudy (xoxolove456):

You wouldn't multiple the 24?